How do I solve the following equation for y?
2/y+4=-8

Answer by praseena(37)


2/y + 4 = -8
subtract 4 from both sides of the equation
==> 2/y = -8 - (-4)
==> 2/y = -8 + 4 { while subtracting a number from a negative number,
we should change the sign of the second number and add
to the first number.
Here we want to subtract -4 from -8
then change the sign of -4 and then add to -8 }


==> 2/y = -4 { here we are adding a positive number and a negative number.
then subtract the smaller number from the bigger number and
put the sign of the bigger.
Here we have -8 and 4.
then first find 8 - 4 = 4 and put the sign of 8, which is
negative.}

now here we want to find out the value of y
so we shopuld bring y to the numerator.
so multiply both sides of the equation by y .
==> y * 2/y = y * -4
==> 2 = -4y
now divide both sides of the equation ny -4
==> 2/-4 = -4y/-4
==> 1/-2 = y
==> y = -1/2

good luck.